OpenStudy (anonymous):

roots of 3x^2-7x+2????

OpenStudy (nathan917):

3x^2 - 27x = 0 x(3x - 27) = 0 x = 0 3x - 27 = 0 3x = 27 x = 27/3 = 9 x = 0 & 9

OpenStudy (anonymous):

how can you do that with the quadratic formula?